If a ball is dropped from the top of a 256-ft building, then the formula
h(t)= 256-16t^2
expresses its height h(t) in feet as a function of the time t in seconds.
1.Find h(2), the height of the ball 2 seconds after it is dropped.
2.Find h(4).

Answer by Alan3354(69443)

If a ball is dropped from the top of a 256-ft building, then the formula
h(t)= 256-16t^2 expresses its height h(t) in feet as a function of the time t in seconds.
1.Find h(2), the height of the ball 2 seconds after it is dropped.
Sub 2 for t
h(2) = 256 - 16*2^2 = 256 - 64
h(2) = 192 ft
------
2.Find h(4).
Do the same as above.